OpenStudy (anonymous):

use Gauss's approach to find the follwoing sums (do not use formulas). a... 1+2+3+4+...+999 b... 1+3+5+7+...+997 the sum of sequance a is... the sum of sequance b is...

OpenStudy (klimenkov):

Write the sum in 2 ways: \[1+2+3+\ldots+999\]\[999+998+997+\ldots+1\]and notice that the sum of 2 numbers that are on the same vertical is the same for any vertical \[(1+999)=(2+998)=\ldots=1000\]Just add them using this hint. If you understood what i had written you will have no problems with getting the sum.